X Company is considering buying a part in 2020 that it has been making for the past several years. A company has offered to supply this part for $15.58 per unit. Budgeted production in 2020 is 56,000 units, and budgeted per-unit production costs are: Materials $6.00 Direct labor [all variable] 4.10 Total overhead 5.00 Total $15.10 $61,600 of X Company's total overhead costs are fixed; $48,664 of the $61,600 are unavoidable even if it buys the part. Also, if X Company buys the part, it can rent the equipment that was used to make the part to another company for $70,000.

3. If X Company continues to make the part instead of buying it, it will save

4. At what production level would X company be indifferent between making and buying the part?
Answer #1

1) Differential analysis

Make Buy
Direct material 336000
Direct labor 229600
Variable overhead 218400
Fixed overhead 12936
Purchase cost 872480
Total relevant cost 796936 872480

it will save = (872480-796936) = 75544

4) Indifferent point

Manufacturing cost = Purchase cost

14X+12936 = 15.58

-1.58X = -12936

X(Indifferent point) = 8187 Units